OpenStudy (anonymous):

In the equation M1V1+M2V2=M1V1'+M2V2' find the formula for the following: V1 , V2 , V1' , V2' Please show/explain how to answer

OpenStudy (anonymous):

if these are just given then for example for V1=(M1V1'+M2V2'-M2V2)/M1

OpenStudy (anonymous):

You may need another equation probably using the coefficient of restitution.

OpenStudy (anonymous):

If v1 and v2 is known, it's easy to find v1' and v2'. (1/2)m1v1^2+(1/2)m2v2^2=(1/2)m1v1'^2+(1/2)m2v2'^2 m1v1^2+m2v2^2=m1v1'^2+m2v2'^2 ...(1) m1(v1^2-v1'^2)=m2(v2'^2-v2^2) ...(2) and the given, m1v1+m2v2=m1v1'+m2v2' ...(a) m1v1-m1v1'=m2v2'-m2v2 m1(v1-v1')=m2(v2'-v2) ...(3) Divide (2) by (3)... v1+v1'=v2+v2' ...(4) m2v1+m2v1'=m2v2+m2v2' (multiplying by m2) ...(5) subtract (5) by (a)... 2m2v2+(m1-m2)v1=(m1+m2)v1' you can find v1' from here. In the same way, multiply (4) by m1, then add to (a)... you can find v2'.